Circular No. 7025

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 J. McClintock and G. Sobczak, Harvard-Smithsonian Center for Astro-

physics; R. Remillard, E. Morgan, and A. Levine, Massachusetts Institute of Technology; C. Bailyn, Yale University; and J. Orosz, Pennsylvania State University, communicate: “XTE J1550–564 was observed with the RXTE PCA during the Sept. 19–20 x-ray flare reported on IAUC 7019. The source intensity during the 3000-s exposure was above 6 Crab. Fourier analysis of the lightcurve from the full energy range of the PCA (effectively 2–30 keV in this case) reveals a high-frequency quasiperiodic oscillation (QPO) at 183.6 Hz, with a width (FWHM) of 46 Hz and an rms amplitude of 0.45 percent of the mean x-ray flux. The statistical significance of the feature is above 8σ. The QPO is not seen during other PCA observations at lower flux. This is the third blackhole candidate known to exhibit x-ray QPOs above 50 Hz, following the 67-Hz QPO in GRS 1915+105 (Morgan et al. 1997, Ap.J. 482, 993) and the 298-Hz QPO in GRO J1655–40 (Remillard

et al. 1998, http://xxx.lanl.gov/abs/astro-ph/?9806049). The QPO

is likely to be an effect of general relativity, with the QPO frequency dependent on both the mass and spin of the accreting black hole, but there are competing models for such oscillations. Meanwhile, recent ASM and PCA observations made between Sept. 25 and Oct. 5 indicate intensities (2–12 keV) in the range of 1.2 to 1.4 Crab.”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), write: “The brightening of the optical counterpart of the Aquila X-1 recurrent transient reported on IAUC 7017 did not develop into an full-blown outburst. CCD photometry obtained with the 1.2-m OHP telescope up until Oct. 3 shows that the source remained at a low intensity level for ∼ 10 days, and some time after Sept. 24 it faded into quiescence, where it now remains. During this brightening, the 18.95-hr modulation (IAUC 6806) was detected with the expected phase and amplitude. We have observed precursor events in 1995, 1996, and 1997 that preceded the main outburst by ∼ 2 months. If this latest event is such a precursor, a normal outburst would be expected by early November.”

1998 October 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green